OpenStudy (anonymous):

use graphing to solve the system of equations 1/2x + 5y=17 and 3x + 2y=18

OpenStudy (anonymous):

1/2x +5y =17 You have to get 5 by it self. You can do that by subtracting 1/2x to get the x to the other side. 5y= -1/2x + 17 Divide 5 now to get Y by it self y= -1/10x + 17/5 Same for this 3x +2y =18 2y= -3x +18 y= -3/2x +9